# Arithmetic Bohr radius and Local Banach space theory

### v1

@@ Abstract.
This article introduces the notion of arithmetic Bohr radius for operator valued pluriharmonic functions on complete Reinhardt domains in $\mathbb{C}^{n}$. Using tools from local Banach space theory, we determine its asymptotic behavior in both finite and infinite dimensions. The framework developed in this article extends the classical Minkowski-space setting to a much broader class of sequence spaces, such as mixed Minkowski, Lorentz, and Orlicz spaces. This generality allows for a unified and systematic investigation of Bohr’s theorem for both holomorphic and pluriharmonic functions.
@@
Pluriharmonic functions, Bohr phenomenon; Power series; complete Reinhardt domain, Minkowski space, Banach sequence space
@@ 1. Introduction and the main results
Over the past three decades, Bohr’s theorem and its applications have become a central topic in the function theory of several complex variables, exhibiting deep connections with fundamental aspects of both the local and global theory of Banach spaces. These rich connections have stimulated extensive research on Bohr’s theorem in various fields of mathematics. For instance, it has been investigated in the contexts of Banach algebras and uniform algebras (see [39, 40]), complex manifolds (see [3, 4]), ordinary and vector-valued Dirichlet series (see [10, 27]), elliptic equations (see [5]), Faber–Green condensers (see [38]), free holomorphic functions (see [42]), vector-valued holomorphic functions (see [30]), local Banach space theory (see [23]), domains of monomial convergence (see [28]), harmonic and pluriharmonic mappings (see [21]), Hardy spaces (see [11]), as well as various multidimensional settings (see [2, 14, 15, 24, 25]). In multidimensional settings, the Bohr radius problem becomes significantly more complicated. Many questions concerning the precise behavior of multidimensional Bohr radii remain open, largely due to the lack of certain essential analytical and geometric tools. Several of these problems are closely connected to fundamental topics in functional analysis, such as the geometry of Banach spaces [13, 31], unconditional basis constants of spaces of polynomials [4], as well as Gordon–Lewis constants, projection constants, and the Banach–Mazur distance [23]. An especially intriguing direction arises when one considers this problem for vector valued functions on complete Reinhardt domains.
@@
$\sup_{z\in r\Omega}\,\sum_{\alpha}\left\lVert U(a_{\alpha})z^{\alpha}\right\rVert^{p}_{Y}\leq\lambda^{p}\,\left\lVert f\right\rVert^{p}_{\Omega,X},$ (1.1)
@@
Let us fix some notations. We write $K(\Omega,p,U):=K_{1}(\Omega,p,U)$. When $U=I:X\rightarrow X$ is the identity operator, we set $K_{\lambda}(\Omega,p,X):=K_{\lambda}(\Omega,p,I)$ and $K(\Omega,p,X):=K_{1}(\Omega,p,X)$. In the scalar valued case, we denote $K_{\lambda}(\Omega,p):=K_{\lambda}(\Omega,p,\mathbb{C})$ and $K(\Omega,p):=K_{1}(\Omega,p)$. The above definition is motivated by Defant et al. [31], extending their result for the case $p=1$ and $\Omega=\mathbb{D}^{n}$ to a more general setting, thus making it possible to the study of Bohr phenomenon for broader classes of domains and functions. With these notations, a classical theorem of Harald Bohr concerning power series states that $K(\mathbb{D},1)=1/3$ (see [16]). One of the most striking features of Bohr’s theorem is that our understanding of the behavior of the constant $K(\Omega,1)$ remains rather limited. A central challenge is to determine the precise value of $K(\Omega,1)$, which is still unknown for $n>1$, even in the classical setting of the polydisc $\mathbb{D}^{n}$, where $\mathbb{D}^{n}:=\{z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}:|z_{j}|<1\,\mbox{for all}\,1\leq j\leq n\}$. To make progress in this direction, much of the existing research has focused on establishing bounds for $K(\Omega,1)$ and investigating its asymptotic behavior on general domains, and in particular well known settings such as the unit ball of Minkowski space and finite-dimensional complex Banach spaces. Denote $B_{\ell^{n}_{q}}:=\left\{z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}:\left\lVert z\right\rVert_{q}:=\left(\sum_{i=1}^{n}|z_{i}|^{q}\right)^{1/q}<1\right\}$, $1\leq q<\infty$ and $B_{\ell^{n}_{\infty}}:=\mathbb{D}^{n}$. For every $1\leq q\leq\infty$ and all $n\in\mathbb{N}$, together the results of Aizenberg, Boas, Dineen, Khavinson, Timoney, Defant, and Frerick from [2, 14, 15, 29] show that there exist constants $C,D>0$ such that
$\frac{1}{C}\,\left(\frac{\log\,n}{n}\right)^{1-\frac{1}{\min\{q,2\}}}\leq K_{n}(B_{\ell^{n}_{q}})\leq D\,\left(\frac{\log\,n}{n}\right)^{1-\frac{1}{\min\{q,2\}}}.$
@@
It is worth mentioning that the constant $K_{\lambda}(\Omega,p,U)$ is not necessarily nonzero for arbitrary Banach spaces $X,Y$, and the operator $U$. We refer to the occurrence of strictly positive constant $K_{\lambda}(\Omega,p,U)$ as the Bohr phenomenon. More precisely, we make the following definition. For given Banach spaces $X,Y$ and the operator $U$ as above, a class of holomorphic functions $f$, defined in a complete Reinhardt domain $\Omega$, is said to exhibit the Bohr phenomenon on $\Omega$ if there exists a universal constant $r=r_{0}\in(0,1]$, called the Bohr radius of $\Omega$ with respect to this class, such that inequality (1.1) is satisfied by every function in the class. Since the Bohr phenomenon does not occur for all classes of holomorphic functions (see [3, 6, 11, 13]), an important problem is to identify and characterize those classes for which the Bohr phenomenon holds. Blasco [13] initiated the systematic study of classes exhibiting the Bohr phenomenon by characterizing them in the classical one-dimensional setting $\Omega=\mathbb{D}$ with $U$ equal to the identity operator. Many thanks to Defant et al. [31], who were the first to address this problem in multidimensional case $\Omega=\mathbb{D}^{n}$, $n>1$, with $p=1$ and for an arbitrary operator $U$. More recently, the problem for general bounded complete Reinhardt domains has been considered by the author in [36], where a framework based on local Banach space theory was developed.
@@
$f(z)=\sum_{m=0}^{\infty}\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{m=1}^{\infty}\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha},$ (1.2)
@@
$\sup_{z\in r\Omega}\,\sum_{m=0}^{\infty}\sum_{|\alpha|=m}(\left\lVert U(a_{\alpha})\right\rVert^{p}_{Y}+\left\lVert U(b_{\alpha})\right\rVert^{p}_{Y})|z^{\alpha}|^{p}\leq\lambda^{p}\,\left\lVert f\right\rVert^{p}_{\Omega,\mathcal{B}(\mathcal{H})}.$ (1.3)
@@
Recently, the author in [36] has shown that the constant $R_{\lambda}(\Omega,p,U)$ is not necessarily nonzero for arbitrary operator $U$. Consequently, in the same paper, the author has established the condition when this constant is non-zero. Moreover, the author investigated asymptotic behaviors of this constant using invariants from local Banach space theory. With these developments in mind, it is natural to ask the following question for the constant $AP_{\lambda}(\Omega,p,U)$.
@@
$AP_{\lambda}(B_{Z},p,U)\geq\frac{\left\lVert Id:Z\rightarrow\ell^{n}_{1}\right\rVert}{n}\,R_{\lambda}(B_{Z},p,U).$ (1.4)
@@
$\lambda^{p}\,|\cos\frac{1}{k}|^{p}+\lambda^{p}\,|\sin\frac{1}{k}|^{p}\,|z_{1}|^{p}\leq\lambda^{p}$ (1.5)
@@
We now impose additional conditions on the initial coefficient, which lead to improved estimates for the arithmetic Bohr radius obtained in Theorem 1.2. More precisely, we consider the following class of functions. Let $\mathcal{PH}^{*}(\Omega,\mathcal{B}(\mathcal{H})):=\{f\in\mathcal{PH}(\Omega,\mathcal{B}(\mathcal{H})):f(0)=\mu\,I,\,\mu\in[0,\left\lVert f\right\rVert_{\Omega,\mathcal{B}(\mathcal{H})}]\}$. Denote $AP^{*}_{\lambda}(\Omega,p,U):=AP_{\lambda}(\mathcal{PH}^{*}(\Omega,\mathcal{B}(\mathcal{H})),p,U)$.
@@
is a norm on $\mathbb{C}^{n}$, and $\Omega$ coincides with the open unit ball of $(\mathbb{C}^{n},p_{\Omega})$. Conversely, if $Z=(\mathbb{C}^{n},||.||)$ is a Banach space whose canonical basis is normalized $1$-unconditional, then its open unit ball $B_{Z}$ is clearly a bounded convex complete Reinhardt domain in $\mathbb{C}^{n}$. Thus, the study of the constant $R_{\lambda}(\Omega,p,U)$ for the unit balls $\Omega=B_{Z}$ of finite dimensional complex Banach spaces $(\mathbb{C}^{n},||.||)$ with the normalized $1$-unconditional canonical bases, is equivalent to studying $R_{\lambda}(\Omega,p,U)$ over bounded complete Reinhardt domains $\Omega\subseteq\mathbb{C}^{n}$.
@@
Let $Z=(\mathbb{C}^{n},||.||)$ be a Banach space such that $\chi(\{e_{k}\}^{n}_{k=1})=1$. Let $X=\mathcal{B}(\mathcal{H})$ be finite dimensional and $\lambda>1$. Then
@@
The correspondence preceding Theorem 1.4 allows Theorem 1.7 to be extended to bounded simply connected convex complete Reinhardt domains. Moreover, combining Theorem 1.6 and Lemma 4.3 yields the asymptotic estimates for all bounded simply connected complete Reinhardt domains, which we state as a corollary.
@@
    Let $Y$ be any complex Hilbert space and $U:\mathcal{B}(\mathcal{H})\rightarrow Y$ be a non-null bounded linear operator with $\left\lVert U\right\rVert<\lambda$. If $Y$ is a $s$-concave Banach lattice with $2\leq s<\infty$ and there is a $1\leq t<s$ such that $U$ is $(s,1)$-summing, then
@@
Let $U:\mathcal{B}(\mathcal{H})\rightarrow Y$ be a non-null bounded liner operator and $\left\lVert U\right\rVert<\lambda$. Then for every $n$ we have
@@
By applying the same arguments as in the pluriharmonic case in Theorem 1.1 and Theorem 1.11, we obtain the relation between the Bohr radius and the arithmetic Bohr radius for any Banach space $X$-valued holomorphic functions.
@@
Theorems 1.12 and 1.13 remain valid for any bounded complete Reinhardt domain, not only for $B_{Z}$.

### v2

@@ Abstract.
This article introduces the notion of arithmetic Bohr radius for operator valued pluriharmonic functions on complete Reinhardt domains in $\mathbb{C}^{n}$. Using tools from local Banach space theory, we determine its asymptotic behavior in both finite and infinite dimensions. Asymptotic estimates for this constant are derived for both convex and non-convex complete Reinhardt domains. The framework developed in this article extends the classical Minkowski-space setting to a much broader class of sequence spaces, such as mixed Minkowski, Lorentz, and Orlicz spaces. Our results also apply to a wide class of Banach sequence spaces, including symmetric and convex Banach spaces. This generality allows for a unified and systematic investigation of Bohr’s theorem for both holomorphic and pluriharmonic functions. As an application of our results, we obtain several consequences extending known results in the scalar valued setting and in the existing literature.
@@
Pluriharmonic functions, Bohr phenomenon; complete Reinhardt domain, Minkowski space, Banach sequence space, symmetric and convex Banach lattice, $p$-summing operator
@@ 1. Introduction and the main results
Over the past three decades, Bohr’s theorem and its applications have become a central topic in the function theory of several complex variables, exhibiting deep connections with fundamental aspects of both the local and global theory of Banach spaces. These rich connections have stimulated extensive research on Bohr’s theorem in various fields of mathematics. For instance, it has been investigated in the contexts of Banach algebras and uniform algebras (see [41, 42]), complex manifolds (see [3, 4]), ordinary and vector-valued Dirichlet series (see [11, 27]), elliptic equations (see [5]), Faber–Green condensers (see [40]), free holomorphic functions (see [44]), vector-valued holomorphic functions (see [30]), local Banach space theory (see [23, 32]), domains of monomial convergence (see [28]), harmonic and pluriharmonic mappings (see [21]), slice regular functions (see [45, 46, 47]), Hardy spaces (see [12]), as well as various multidimensional settings (see [2, 15, 16, 24, 25]).
The investigation of Bohr’s theorem for slice regular functions has recently attracted considerable attention in hypercomplex analysis. The pioneering work of Rocchetta, Gentili, and Sarfatti [45] in $2012$ established the Bohr’s theorem for slice regular functions, providing a fundamental extension of the classical Bohr phenomenon to the setting of noncommutative function theory. Their results opened new directions in the study of analytic properties of quaternionic and related function spaces. Subsequently, Xu [46] in $2021$ made significant contributions by introducing the notion of the generalized Bohr radius for slice regular functions over quaternions, thereby refining and extending earlier results. In the same year, Xu [47] further extended the Bohr’s theorem to slice regular functions over octonions, addressing additional challenges arising from nonassociativity. These developments have considerably enriched the theory and stimulated further research on Bohr’s theorem in hypercomplex frameworks.
In multidimensional settings, the Bohr radius problem becomes significantly more complicated. Many questions concerning the precise behavior of multidimensional Bohr radii remain open, largely due to the lack of certain essential analytical and geometric tools. Several of these problems are closely connected to fundamental topics in functional analysis, such as the geometry of Banach spaces [14, 31], unconditional basis constants of spaces of polynomials [4], as well as Gordon–Lewis constants, projection constants, and the Banach–Mazur distance [23]. An especially intriguing direction arises when one considers this problem for vector valued functions on complete Reinhardt domains.
@@
(1.1) $\sup_{z\in r\Omega}\,\sum_{\alpha}\left\lVert U(a_{\alpha})z^{\alpha}\right\rVert^{p}_{Y}\leq\lambda^{p}\,\left\lVert f\right\rVert^{p}_{\Omega,X},$
@@
Let us fix some notations. We write $K(\Omega,p,U):=K_{1}(\Omega,p,U)$. When $U=I:X\rightarrow X$ is the identity operator, we set $K_{\lambda}(\Omega,p,X):=K_{\lambda}(\Omega,p,I)$ and $K(\Omega,p,X):=K_{1}(\Omega,p,X)$. In the scalar valued case, we denote $K_{\lambda}(\Omega,p):=K_{\lambda}(\Omega,p,\mathbb{C})$ and $K(\Omega,p):=K_{1}(\Omega,p)$.
@@
The above definition is motivated by Defant et al. [31], extending their result for the case $p=1$ and $\Omega=\mathbb{D}^{n}$ to a more general setting, thus making it possible to the study of Bohr phenomenon for broader classes of domains and functions. With the above notations, a classical theorem of Harald Bohr concerning power series states that $K(\mathbb{D},1)=1/3$ (see [17]). One of the most striking features of Bohr’s theorem is that our understanding of the behavior of the constant $K(\Omega,1)$ remains rather limited. A central challenge is to determine the precise value of $K(\Omega,1)$, which is still unknown for $n>1$, even in the classical setting of the polydisc $\mathbb{D}^{n}$, where $\mathbb{D}^{n}:=\{z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}:|z_{j}|<1\,\mbox{for all}\,1\leq j\leq n\}$. To make progress in this direction, much of the existing research has focused on establishing bounds for $K(\Omega,1)$ and investigating its asymptotic behavior on general domains, and in particular well known settings such as the unit ball of Minkowski space and finite-dimensional complex Banach spaces.
Denote $B_{\ell^{n}_{q}}:=\left\{z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}:\left\lVert z\right\rVert_{q}:=\left(\sum_{i=1}^{n}|z_{i}|^{q}\right)^{1/q}<1\right\}$, $1\leq q<\infty$ and $B_{\ell^{n}_{\infty}}:=\mathbb{D}^{n}$. For every $1\leq q\leq\infty$ and all $n\in\mathbb{N}$, together the results of Aizenberg, Boas, Dineen, Khavinson, Timoney, Defant, and Frerick from [2, 15, 16, 29] show that there exist constants $C,D>0$ such that
$\frac{1}{C}\,\left(\frac{\log\,n}{n}\right)^{1-\frac{1}{\min\{q,2\}}}\leq K(B_{\ell^{n}_{q}},1)\leq D\,\left(\frac{\log\,n}{n}\right)^{1-\frac{1}{\min\{q,2\}}}.$
Boas and Khavinson [15] showed that the constant $K(\mathbb{D}^{n},1)$ does not exceed $1/3$, and moreover that the sequence $\{K(\mathbb{D}^{n},1)\}_{n=1}^{\infty}$ is decreasing with limit zero as $n\rightarrow\infty$. In the bidisc case, their work yields the bounds $0.23570226\leq K(\mathbb{D}^{2},1)\leq 1/3=0.33333$. Considerable progress has been made only recently. Knese [38] established in $2025$ that $K(\mathbb{D}^{2},1)>0.3006$, while a subsequent result in $2026$ of Baran, Pikul, Woerdeman, and Wojtylak [9] showed that $K(\mathbb{D}^{2},1)<0.3177$. These advances in particular ensure that $K(\mathbb{D}^{n},1)<1/3$ for $n\geq 2$.
@@
It is worth mentioning that the constant $K_{\lambda}(\Omega,p,U)$ is not necessarily nonzero for arbitrary Banach spaces $X,Y$, and the operator $U$. We refer to the occurrence of strictly positive constant $K_{\lambda}(\Omega,p,U)$ as the Bohr phenomenon. More precisely, we make the following definition. For given Banach spaces $X,Y$ and the operator $U$ as above, a class of holomorphic functions $f$, defined in a complete Reinhardt domain $\Omega$, is said to exhibit the Bohr phenomenon on $\Omega$ if there exists a universal constant $r=r_{0}\in(0,1]$, called the Bohr radius of $\Omega$ with respect to this class, such that inequality (1.1) is satisfied by every function in the class. Since the Bohr phenomenon does not occur for all classes of holomorphic functions (see [3, 6, 12, 14]), an important problem is to identify and characterize those classes for which the Bohr phenomenon holds. Blasco [14] initiated the systematic study of classes exhibiting the Bohr phenomenon by characterizing them in the classical one-dimensional setting $\Omega=\mathbb{D}$ with $U$ equal to the identity operator. Many thanks to Defant, Maestre, and Schwarting [31], who were the first to address this problem in multidimensional case $\Omega=\mathbb{D}^{n}$, $n>1$, with $p=1$ and for an arbitrary operator $U$. More recently, the problem for general bounded complete Reinhardt domains has been considered by the present author in [37], where a framework based on local Banach space theory was developed.
@@
(1.2) $f(z)=\sum_{m=0}^{\infty}\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{m=1}^{\infty}\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha},$
@@
(1.3) $\sup_{z\in r\Omega}\,\sum_{m=0}^{\infty}\sum_{|\alpha|=m}(\left\lVert U(a_{\alpha})\right\rVert^{p}_{Y}+\left\lVert U(b_{\alpha})\right\rVert^{p}_{Y})|z^{\alpha}|^{p}\leq\lambda^{p}\,\left\lVert f\right\rVert^{p}_{\Omega,\mathcal{B}(\mathcal{H})}.$
@@
Recently, the present author in [37] has shown that the constant $R_{\lambda}(\Omega,p,U)$ is not necessarily nonzero for arbitrary operator $U$. Consequently, in the same paper, the author has established the condition when this constant is non-zero. Moreover, the author investigated asymptotic behaviors of this constant using invariants from local Banach space theory. With these developments in mind, it is natural to ask the following question for the constant $AP_{\lambda}(\Omega,p,U)$.
@@
(1.4) $AP_{\lambda}(B_{Z},p,U)\geq\frac{\left\lVert Id:Z\rightarrow\ell^{n}_{1}\right\rVert}{n}\,R_{\lambda}(B_{Z},p,U).$
@@
Applying this corollary to the case $\Omega=B_{\ell^{n}_{q}}$, the Minkowski space, yields the following result.
Corollary 1.2.
Let $Y$ be any complex Banach space and $U:\mathcal{B}(\mathcal{H})\rightarrow Y$ a bounded liner operator with $\left\lVert U\right\rVert\leq\lambda$. For each $1\leq p<\infty$ and $1\leq q\leq\infty$,
$AP_{\lambda}(B_{\ell^{n}_{q}},p,U)\geq D\,\begin{cases}n^{-\frac{1}{q}}&\text{for $q\leq p$},\\[5.69054pt] n^{-\frac{1}{p}}&\text{for $q>p$},\end{cases}$
where $D$ as in Theorem 1.1.
@@
(1.5) $\lambda^{p}\,|\cos\frac{1}{k}|^{p}+\lambda^{p}\,|\sin\frac{1}{k}|^{p}\,|z_{1}|^{p}\leq\lambda^{p}$
@@
We now impose additional conditions on the initial coefficient, which lead to improve estimates for the arithmetic Bohr radius obtained in Theorem 1.2. More precisely, we consider the following class of functions. Let $\mathcal{PH}^{*}(\Omega,\mathcal{B}(\mathcal{H})):=\{f\in\mathcal{PH}(\Omega,\mathcal{B}(\mathcal{H})):f(0)=\mu\,I,\,\mu\in[0,\left\lVert f\right\rVert_{\Omega,\mathcal{B}(\mathcal{H})}]\}$. Denote $AP^{*}_{\lambda}(\Omega,p,U):=AP_{\lambda}(\mathcal{PH}^{*}(\Omega,\mathcal{B}(\mathcal{H})),p,U)$.
@@
is a norm on $\mathbb{C}^{n}$, and $\Omega$ coincides with the open unit ball of $(\mathbb{C}^{n},p_{\Omega})$. Conversely, if $Z=(\mathbb{C}^{n},||.||)$ is a Banach space whose canonical basis is normalized $1$-unconditional, then its open unit ball $B_{Z}$ is clearly a bounded convex complete Reinhardt domain in $\mathbb{C}^{n}$. Thus, the study of the constant $R_{\lambda}(\Omega,p,U)$ for the unit balls $\Omega=B_{Z}$ of finite dimensional complex Banach spaces $(\mathbb{C}^{n},||.||)$ with the normalized $1$-unconditional canonical bases, is equivalent to studying $R_{\lambda}(\Omega,p,U)$ over bounded convex complete Reinhardt domains $\Omega\subseteq\mathbb{C}^{n}$.
@@
Let $X=\mathcal{B}(\mathcal{H})$ be finite dimensional and $\lambda>1$. Then, for $1\leq q\leq\infty$,
@@
The correspondence preceding Theorem 1.4 allows Theorem 1.7 to be extended to bounded simply connected convex complete Reinhardt domains. Moreover, combining Theorem 1.6 with Lemma 4.3 yields the asymptotic estimates for all bounded simply connected complete Reinhardt domains, not necessarily convex. We record this result as a corollary.
@@
    Let $Y$ be any complex Banach space and $U:\mathcal{B}(\mathcal{H})\rightarrow Y$ be a non-null bounded linear operator with $\left\lVert U\right\rVert<\lambda$. If $Y$ is a $s$-concave Banach lattice with $2\leq s<\infty$ and there is a $1\leq t<s$ such that $U$ is $(s,1)$-summing, then
@@
Let $U:\mathcal{B}(\mathcal{H})\rightarrow Y$ be a non-null bounded liner operator and $\left\lVert U\right\rVert<\lambda$. Then for every $n\in\mathbb{N}$, we have
@@
It worth mentioning that Corollary 1.8 is actually pluriharmonic analogue of [27, Theorem 3.1].
In view of the corollaries 1.6, 1.7, and 1.8, it is evident that to estimate $AP_{\lambda}(\Omega,p,U)$ it is enough to estimate $R_{\lambda}(\mathbb{D},p,U)$. To do this, we prove the following result in the case $U$ is the identity operator on $\mathbb{C}$ and $p=1$.
Lemma 1.1.
If $U=I:\mathbb{C}\rightarrow\mathbb{C}$ and $1<\lambda$, then
$R_{\lambda}(\mathbb{D},1,\mathbb{C})\geq\frac{(\lambda-1)\pi}{4+(\lambda-1)\pi}.$
Now, by making use of Lemma 1.1 and the corollaries 1.6, 1.7, and 1.8, we deduce the following important estimates, which we record as a corollary.
Corollary 1.9.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded simply connected complete Reinhardt domain. Let $U$ be the identity operator on $\mathbb{C}$ and $1<\lambda$. Then for every $n\in\mathbb{N}$, we have
$\frac{(\lambda-1)\pi}{(4+(\lambda-1)\pi)\,n}\leq AP_{\lambda}(\Omega,1,\mathbb{C}).$
This corollary actually provides a universal lower bound of the arithmetic Bohr radius for complex valued pluriharmonic functions on any bounded simply connected complete Reinhardt domain $\Omega\subset\mathbb{C}^{n}$.
On the other hand, by applying the same arguments as in the pluriharmonic case in Theorem 1.1, Theorem 1.11, and Corollary 1.6, we obtain the relation between the Bohr radius and the arithmetic Bohr radius for any Banach space $X$-valued holomorphic functions.
@@
Theorems 1.12 and 1.13 remain valid for any bounded complete Reinhardt domain, not only for $B_{Z}$. Moreover, we recover [27, Lemma 4.3] from Theorem 1.12 by choosing $U=I:\mathbb{C}\rightarrow\mathbb{C}$ and $p=1$.
In particular, Theorem 1.13 yields the following result.
Corollary 1.10.
Let $U:X\rightarrow Y$ be a non-null bounded liner operator and $\left\lVert U\right\rVert<\lambda$. Then for every $n\in\mathbb{N}$ we have
$A_{\lambda}(B_{\ell^{n}_{1}},1,U)=\frac{K_{\lambda}(\mathbb{D},1,U)}{n}.$
It is interesting to note that [27, Theorem 3.1] follows from Corollary 1.10 by choosing $U$ the identity operator on $\mathbb{C}$.